If 3.0 mol CO reacts with 1.0 mol 02



1
Expert's answer
2022-04-15T15:21:01-0400

2CO + O2 = 2CO2


CO excess reagent

Olimiting reagent

n(CO2) = 2n(O2) = 2*1 mol = 2 mol

V(CO2) = 2*22.4= 44.8L

m(CO2) = 2 * 44 = 88 g